A sample of 60 Grade 11 student’s ages was obtained to estimate the mean age of all Grade 11 students. X=16.7 years and the population variance is 16.

  1. Find the 90% confidence interval for ?


  1. Find the 95% confidence interval for ?


  1. What conclusions can you make based on each estimate
1
Expert's answer
2021-06-15T09:25:01-0400

1.1

The critical value for "\alpha=0.1" is "z_c=z_{1-\alpha/2}=1.6449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(16.7-1.6449\times\dfrac{4}{\sqrt{60}}, 16.7+1.6449\times\dfrac{4}{\sqrt{60}})"

"\approx(15.8506, 17.5494)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "15.8506<\mu<17.5494," which indicates that we are 90% confident that the true population mean "\mu"

is contained by the interval "(15.8506, 17.5494)."


1.2

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(16.7-1.96\times\dfrac{4}{\sqrt{60}}, 16.7+1.96\times\dfrac{4}{\sqrt{60}})"

"\approx(15.6879, 17.7121)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "15.6879<\mu<17.7121," which indicates that we are 95% confident that the true population mean "\mu"

is contained by the interval "(15.6879, 17.7121)."


1.3

The width of the confidence interval will be larger when the confidence level is higher (because you can have greater confidence when you are less precise).